# YOLO11 and Vision Transformers based 3D Pose Estimation of Immature Green Fruits in Commercial Apple Orchards for Robotic Thinning

### v1

@@ Abstract
In this study, a robust method for 3D pose estimation of immature green apples (fruitlets) in commercial orchards was developed, utilizing the YOLO11 object pose detection model alongside Vision Transformers (ViT) for depth estimation. For object detection and pose estimation, performance comparisons of YOLO11 (YOLO11n, YOLO11s, YOLO11m, YOLO11l and YOLO11x) and YOLOv8 (YOLOv8n, YOLOv8s, YOLOv8m, YOLOv8l and YOLOv8x) were made under identical hyperparameter settings among the all configurations. Likewise, for RGB to RGB-D mapping, Dense Prediction Transformer (DPT) and Depth Anything V2 were investigated. It was observed that YOLO11n surpassed all configurations of YOLO11 and YOLOv8 in terms of box precision and pose precision, achieving scores of 0.91 and 0.915, respectively. Conversely, YOLOv8n exhibited the highest box and pose recall scores of 0.905 and 0.925, respectively. Regarding the mean average precision at 50% intersection over union (mAP@50), YOLO11s led all configurations with a box mAP@50 score of 0.94, while YOLOv8n achieved the highest pose mAP@50 score of 0.96. In terms of image processing speed, YOLO11n outperformed all configurations with an impressive inference speed of 2.7 ms, significantly faster than the quickest YOLOv8 configuration, YOLOv8n, which processed images in 7.8 ms. This demonstrates a substantial improvement in inference speed over previous iterations, particularly evident when comparing YOLO11n and YOLOv8n. Subsequent integration of ViTs for the green fruit’s pose depth estimation revealed that Depth Anything V2 outperformed Dense Prediction Transformer in 3D pose length validation, achieving the lowest Root Mean Square Error (RMSE) of 1.52 and Mean Absolute Error (MAE) of 1.28, demonstrating exceptional precision in estimating immature green fruit lengths. Following this, the DPT showed notable accuracy improvements with a RMSE of 3.29 and an MAE of 2.62. In contrast, measurements derived from Intel RealSense point clouds exhibited the highest discrepancies from the ground truth, with a RMSE of 9.98 and an MAE of 7.74. These findings emphasize the effectiveness of YOLO11 in detecting and estimating the pose of immature green fruits, illustrating how Vision Transformers like Depth Anything V2 adeptly convert RGB images into RGB-D data, thus enhancing the precision and computational requirement of 3D pose estimations for future robotic thinning applications in commercial orchards.
Keywords YOLO11  $\cdot$ Vision Transformers  $\cdot$ Depth Anything  $\cdot$ Dense Prediction Transformer  $\cdot$ YOLO
@@ 1 Introduction
In United States (US) agriculture, labor shortage is one of the biggest problem Richards et al. (2024); Karan and Asgari (2021) , and automation and robotics have been instrumental in reducing labor efforts and enhancing crop production Vasconez et al. (2019); Grieve et al. (2019). For example, precision planting systems automate seed placement with exact spacing and depth Khadatkar et al. (2021), optimizing growth conditions and reducing the need for manual labor. Likewise, robotic weed control systems identify and eliminate weeds without human intervention, using advanced sensors and precise application of herbicides or mechanical removal techniques Wu et al. (2020); Upadhyay et al. (2024). Furthermore, in harvesting, autonomous combine harvesters and fruit-picking robots increase efficiency by performing tasks faster and with greater accuracy than human workers Xiao et al. (2024); Fei and Vougioukas (2021). Similarly, automated irrigation systems use weather forecasts and soil sensors to regulate water delivery, minimizing wastage and labor associated with manual watering Gutiérrez et al. (2013).
However, the commercial production of specialty tree fruit crops remains heavily reliant on extensive manpower because of paramount need of precision and care for crop-load management operations such as tree pruning, fruit thinning, and harvesting Kaur et al. (2023); Rutledge and Mérel (2023). For example, apple (Malus domestica) orchards require significant labor inputs for pruning to shape the trees and improve sun exposure to the fruits, which is essential for enhancing fruit quality and yield Musacchi and Serra (2018). Similarly, the peach (Prunus persica) industry depends on manual thinning to regulate fruit size and prevent branch damage due to overbearing Iglesias et al. (2023); Costa and Botton (2022). Likewise, cherries (Prunus avium), known for their delicate nature, necessitate careful hand-picking to avoid damage to both the fruit and the trees Karkee et al. (2018), a task that machines are currently unable to perform with sufficient gentleness Zhou et al. (2016). Additionally, in the cultivation of pears (Pyrus communis), precise manual thinning is critical to avoid alternation in bearing years and to ensure consistent fruit quality Bound (2022). Additionally, plums (Prunus domestica) also require extensive manual labor for thinning and pruning to maintain fruit quality and manage tree health Seehuber et al. (2011).
@@
Among these crop-load management operations, immature green fruit thinning, a centuries-old practice is essential in the commercial cultivation of tree fruit crops for optimization of fruit size and quality by reducing competition for water, sunlight, and nutrients among fruits Dennis Jr (2000); Costa et al. (2018). This practice is particularly crucial for apple production, which ranks second in consumption worldwide among major fruits O’Rourke (2021) and the third most consumed fruit in the United States Statista (2022). The challenges of manual thinning, combined with persistent labor shortages and the increasing costs associated with securing skilled labor, underscore the urgent need for innovative solutions Assirelli et al. (2018); Lordan et al. (2018). Each year, over a hundred thousand of worker come to the US for farm labor Holmes and Ramirez-Lopez (2023); Weiler (2022). Example of such farm workers are shown in Figure 1, which was taken in the commercial apple orchard in Prosser, Washington State, USA. The Figure 1 visually encapsulates the labor-intensive nature of fruit thinning in commercial orchards, highlighting the traditional methods still in use. Figure 1a illustrates the painstaking process of manual fruitlet thinning performed by a worker in a commercial ’Scilate’ apple orchard in Prosser, Washington State, USA, during the year 2024. The image captures a laborer navigating between trees with an aluminum ladder, which he uses to access fruit clusters up to 14 feet high. This method not only demands physical agility to climb and reposition the ladder continuously but also poses safety risks inherent in working at heights.
In contrast, Figure 1b depicts an alternative approach where multiple workers operate from a mechanized platform that navigates the rows of the orchard. While this method reduces the need to climb, it still requires significant manual effort. Workers must bend and stretch extensively to reach fruitlets in dense clusters, as demonstrated in Figure 1c. In manual fruit thinning, workers are shown actively managing high-density areas of immature green fruits, a necessary task given the overproduction tendencies of fruit trees. This physically demanding work not only entails high labor costs but also poses significant health risks to workers, particularly related to spinal injuries and other musculoskeletal disorders. Studies have highlighted that the repetitive and strenuous postures required in manual thinning, such as excessive back bending and upper arm elevation, can lead to severe physical strain Thamsuwan et al. (2020). Specifically, while platforms can reduce upper arm strain compared to ladders, they do not significantly alleviate the stress on workers’ backs, who must still engage in substantial torso flexion. Furthermore, the agricultural sector is notorious for the prevalence of upper limb injuries, with a systematic review revealing that a substantial number of these injuries occur during manual tasks like harvesting, often leading to long-term disabilities and economic consequences Mucci et al. (2020).
Despite the advent of some mechanization, the reliance on manual labor in fruit thinning continues due to the precision required in these tasks, underscoring an urgent need for more innovative, automated solutions. Such advancements would not only enhance efficiency but also crucially reduce the physical strain on workers, potentially decreasing the incidence of work-related injuries in commercial fruit production. The current state of labor in agriculture calls for a shift towards systems that prioritize worker health and safety by integrating technology capable of replicating the delicate nature of manual thinning without the associated physical risks.
The primary goal of this study is to address the challenges associated with manual thinning in tree fruit agriculture, especially during the immature stages of the fruit when they are all green and similar in appearance. We employ advanced machine vision techniques to facilitate automation in managing orchards. Using the latest variant, YOLO11 (You Only Look Once, version 11), a sophisticated deep learning algorithm, our study aims to develop a system that can accurately detect and determine the position and orientation, or ’pose’, of immature fruitlets in standard RGB (red, green, blue) image space. Further, we enhance the precision of our pose estimations by integrating the Depth Anything model, a tool that performs monocular depth estimation from RGB images. By comparing these estimations to actual ground truth measurements in RGB-D (RGB + Depth) space, we validate the effectiveness of our approach, setting the stage for subsequent automation efforts in orchard management.
The specific objectives of this study can be summarized into following points:
  • Data Acquisition: To deploy a robotic platform to collect high-resolution RGB images of immature fruitlets within a commercial orchard environment for assembling a comprehensive dataset.
  • Deep Learning based Detection and Pose Estimation: To train and assess the performance of the YOLO11 and YOLOv8 models under similar hyperparameter settings to detect and estimate the pose of immature green fruits. Implement the best-performing model to estimate the pose of fruitlets from RGB images.
  • RGB to RGB-D Mapping Using Vision Transformers: To convert the RGB images into 3D point clouds to enhance the pose estimations provided by the best-performing deep learning model using Vision Transformers: Dense Prediction Transformer (DPT) and Depth Anything V2.
  • Field Level Validation: To validate these depth estimations against field-level measurements taken in commercial orchards, comparing the major axis values of immature green apple fruits, collected with vernier caliper, to their respective point cloud data.
@@
Figure 1: Illustrates various manual thinning methods in a commercial ’Scilate’ apple orchard in Prosser, Washington State, USA, during 2024. Subfigure (a) depicts a worker using an aluminum ladder to access high fruit clusters, demonstrating the physical demands and safety risks of traditional methods. Subfigure (b) shows workers on a mechanized platform, highlighting reduced climbing but still extensive manual effort required in dense clusters. Subfigure (c) captures workers managing high-density fruit areas, emphasizing the labor-intensive nature of the task.
2 Related Works
@@ 2.1 Deep Learning-based Fruit Detection and Pose Estimation:
Traditional methods for apple detection, crucial for robotic thinning, have predominantly focused on mature fruit suitable for harvesting, utilizing basic image processing techniques. These methods, however, often fall short in complex natural environments where young fruits exhibit colors and textures similar to foliage, complicating their detection Kataoka et al. (1998); Xuan et al. (2020). Recent advancements in machine learning and deep learning have introduced more sophisticated models, such as YOLOv3 and its variations, which significantly enhance detection capabilities by learning from diverse datasets without explicit programming Liakos et al. (2018); Xuan et al. (2020). Despite these technological advances, few studies have targeted the detection systems necessary for robotic thinning of early-stage apples. The specific challenges during this growth stage include the obscurity and overlap of fruits by leaves and branches, as well as variable lighting conditions, which require a robust detection system capable of operating under such complexities Linker et al. (2012); Sun et al. (2019). Recent efforts have shown promising results using deep learning techniques for young apple detection, yet the need for higher accuracy and faster processing speeds remains Xia et al. (2018); Tian et al. (2019); Huang et al. (2021).
Likewise, Eizentals et al. Eizentals and Oka (2016) introduced a 3D pose estimation method for green peppers using Lidar-based model matching. Although effective under controlled conditions, this method’s dependency on precise laser range findings and its moderate error rates highlight the need for more accurate technologies in cluttered and variable outdoor environments. Lin et al. (2019) explored guava detection and pose estimation using a low-cost RGB-D sensor, achieving high precision and recall. However, the pose error and the 0.565 seconds per fruit processing time underscore potential limitations in real-time applications, indicating a demand for faster and more precise pose estimation techniques. Yin et al. (2021) utilized binocular imagery and deep neural networks for grape cluster-harvesting robots. Their approach demonstrated high accuracy in fruit detection and pose estimation. Yet, the 1.7 seconds processing time per grape and the reliance on complex setup for binocular imagery point to a need for simpler and more efficient methods. Sun et al. (2023) described a citrus pose estimation method using a single RGB image, focusing on fruit rotation vector prediction. While innovative, the method’s reliance on specific image annotations and the moderate harvesting success rate suggest room for improvement in pose accuracy and operational efficiency. Li et al. (2018) developed a pose estimation method for sweet peppers by detecting symmetry axes in point clouds. Despite its precision, the method’s performance varies significantly with the physical state of the peduncle, highlighting a limitation in consistency and the potential for errors in uncontrolled environments.
Additionally, Kim et al. Kim et al. (2023) developed a 2D pose estimation method for tomatoes, emphasizing pedicel detection for robotic harvesting. While this approach effectively grasps the poses of tomato-pedicel pairs, the reliance on visible keypoints may falter under occlusion or varied postures, indicating the need for more robust methods in cluttered environments. Giefer et al. (2019) employed deep learning for apple pose estimation in logistic centers, achieving high accuracy in rotation regression. However, the method’s reliance on single-perspective imaging may limit its applicability in more complex orchard settings, where multi-angle approaches could enhance detection accuracy. Jang and Hwang (2024) utilized the association between tomato bodies and sepals for pose estimation, achieving promising accuracy. Nonetheless, the method’s dependency on high-quality association metrics like IoU scores could be problematic in less controlled field conditions, suggesting a gap in versatility and robustness. Guo et al. (2020) presented a 6D pose estimation strategy using point cloud registration for various fruits. While effective, the need for extensive pre-processing and 3D model matching might not scale efficiently across diverse fruit types without significant computational resources. Luo et al. (2022) combined point cloud segmentation and geometric analysis for in-field grape cluster pose estimation. Though the method is functional, the reported error angles in pose estimation and the time required per cluster highlight the ongoing challenges in achieving real-time, precision-oriented harvesting solutions.
@@ 2.2 Vision Transformers Application in Agriculture
Vision Transformers (ViT) have emerged as powerful alternatives to convolutional neural networks for various image analysis tasks, leveraging the self-attention mechanism to process images as sequences of patches. The Dense Prediction Transformer (DPT), an adaptation of ViT, is designed for dense prediction tasks such as semantic segmentation and depth estimation. DPT modifies the traditional transformer architecture to handle high-resolution inputs directly, maintaining spatial hierarchies and enabling precise localization capabilities crucial for detailed image understanding tasks.
In the emerging field of precision agriculture, vision transformers are increasingly recognized for their efficacy in addressing complex challenges such as plant disease classification. Yet, despite their burgeoning utility, the application of vision transformers within agricultural contexts, especially beyond disease identification, remains relatively sparse. Salamai et al. (2023) introduced a lesion-aware visual transformer specifically designed for the detection of paddy diseases, marking a significant advance with its multi-scale learning and novel lesion localization capabilities that guide disease detection. This study showcased an exceptional model performance with accuracy and F1-score averaging 98.74% and 98.18%, respectively, underscoring its potential in precision agriculture. Thai et al. (2023) developed the FormerLeaf, a transformer-based model that employs the Least Important Attention Pruning (LeIAP) algorithm and sparse matrix-matrix multiplication (SPMM) to enhance the detection of leaf diseases. The innovations presented not only reduced training time by 10% but also increased accuracy, demonstrating the transformer’s ability to adapt and perform efficiently in agricultural settings. Thakur et al. (2021) presented the PlantViT, a hybrid model that combines CNNs with a Vision Transformer to diagnose plant diseases from leaf images. This model, which leverages multi-head attention modules, achieved notable accuracy on extensive datasets, emphasizing the transformer’s robustness and generalizability in diverse agricultural environments. Li et al. (2022) proposed an automatic pest identification method using a Vision Transformer, trained on an enhanced dataset to prevent overfitting and improve accuracy. The results indicated that the ViT network outperformed traditional CNN-based methods, highlighting the transformer’s precision in classifying plant diseases and pests. Lastly, Pacal (2024) exploited an advanced vision transformer model to identify diseases in maize leaves, achieving a groundbreaking accuracy rate of 99.24%. By integrating a MaxViT model with novel architectural adjustments, this study set a new benchmark for deep learning applications in agriculture, proving that vision transformers can significantly enhance crop productivity and sustainability.
@@
Beyond plant disease detection, the utilization of vision transformers in agriculture extends to innovative applications such as crop yield prediction and multi-object tracking, as demonstrated by two notable studies. Lin et al. (2023) introduced the Multi-Modal Spatial-Temporal Vision Transformer (MMST-ViT), a pioneering model designed for precise crop yield prediction. This model uniquely integrates visual remote sensing data with meteorological inputs to assess the impact of both short-term weather variations and long-term climate change on crop yields. The MMST-ViT consists of three specialized components: a Multi-Modal Transformer for processing diverse data types, a Spatial Transformer for analyzing spatial dependencies across regions, and a Temporal Transformer for capturing long-range temporal dynamics. The novel multi-modal contrastive learning technique employed allows for effective pre-training without extensive human supervision. This approach has been validated across over 200 counties in the United States, where it consistently outperformed existing models on several performance metrics. Similarly, Hernandez and Medeiros (2024) developed a Multi-Object Tracking (MOT) framework tailored for agricultural environments using a Vision Transformer, enhancing the robustness and accuracy of robotic tracking applications. Their framework utilizes the local feature matching transformer (LoFTR), to establish spatial associations among plants within cluttered fields, significantly improving tracking performance. Tested on the LettuceMOT and AppleMOT benchmark datasets, their method showed substantial enhancements in tracking accuracy, with improvements up to 44% over existing solutions. This underscores the transformative potential of Vision Transformers in managing the spatial dynamics of plants, facilitating advancements in robotic agriculture.
ViTs have demonstrated significant promise across various domains, including precision agriculture, primarily focusing on plant disease detection. The groundbreaking contributions by Salamai et al. (2023), Thai et al. (2023), Thakur et al. (2021), Li et al. (2022), and Pacal (2024) have showcased the capabilities of ViTs in enhancing classification accuracy and processing efficiency in this specific area. However, the broader application of these models in agriculture, particularly in tasks such as crop yield prediction and multi-object tracking, is still in its nascent stages. The introduction of the MMST-ViT model by Lin et al. (2023) and the multi-object tracking framework by Hernandez and Medeiros (2024) are pioneering efforts that expand the use of ViTs beyond disease detection.
These studies illustrate the adaptability of ViTs to diverse agricultural needs. Despite these advances, the integration of ViTs into wider agricultural applications, such as real-time automated systems for crop management and harvesting robotics, remains limited. The agricultural sector faces challenges in integrating ViTs with existing technologies, addressing compatibility with various sensor types, and meeting real-time processing demands. To fully leverage the potential of ViTs in agriculture, further research is essential. This research should focus on overcoming the current limitations and expanding the application of ViTs to more complex agricultural environments. Specifically, there is a compelling opportunity to explore the use of ViTs in environments such as apple orchards and other specialty crops, which undergo similar phases of management, pruning, and thinning throughout their growth cycles. These environments, where crops present similar challenges to those faced in apple orchard management, could greatly benefit from the advanced image analysis capabilities of ViTs. Such studies could pave the way for comprehensive, efficient, and adaptable agricultural practices, maximizing the transformative potential of Vision Transformers in agriculture.

### v2

@@ Abstract
Manual thinning of immature green apples in commercial orchards is labor-intensive and costly, necessitating the development of automated solutions to address labor shortages and improve efficiency. This study developed a robust method for 3D pose estimation of immature green apples (fruitlets) in commercial orchards using YOLO11 object detection model and Vision Transformers for depth estimation. The results revealed that YOLO11n outperformed other configurations in box and pose precision (0.91 and 0.92, respectively) and demonstrated the fastest inference speed of 2.7 ms per image. YOLO11s achieved the highest box mAP@50 score of 0.94, while YOLOv8n reached the highest pose mAP@50 score of 0.96. For depth estimation, Depth Anything V2 showed superior performance with the lowest RMSE of 1.52 and MAE of 1.28 in 3D pose length validation. These findings highlight the potential of integrating YOLO11 and Vision Transformers, particularly Depth Anything V2, for accurate and efficient 3D pose estimation of immature green fruits, paving the way for future robotic thinning applications in commercial orchards, which is expected to reduce the labor demands in orchard operations.
keywords:
Greenfruit Pose Estimation , YOLO11 pose detection , deep learning in agriculture , AI in Agriculture , Agricultural Automation , Machine-Vision
journal: Computers and Electronics in Agriculture
[…]
[1]organization=Biological & Environmental Engineering, addressline=Cornell University, city=Ithaca, postcode=14850, state=NY, country=USA
[…]
[2]organization=Center for Precision and Automated Agricultural Systems, Department of Biological Systems Engineering, addressline=Washington State University, state=WA, country=USA
@@ 1 Introduction
Labor shortages continue to present significant challenges in U.S. agriculture, prompting the need for advancing automation and robotics solutions to minimize the dependence on manual labor and enhance crop production Richards et al. (2024); Karan and Asgari (2021). Technologies such as precision planting systems automate seed placement with exact spacing and depth Khadatkar et al. (2021), optimizing growth conditions and reducing the need for manual labor. Similarly, robotic weed control systems commercialized in recent years use advanced sensors and precise application/weeding techniques to identify and eliminate weeds without human intervention Wu et al. (2020); Upadhyay et al. (2024). Researchers and technology companies are also developing fruit-picking robots with capabilities to harvest crops such as apples and citrus as fast or faster than human pickers and with greater accuracy Xiao et al. (2024); Fei and Vougioukas (2021).
Despite advancements in automation and robotics for various types of crops, field operations in tree fruit crops such as apples (Malus domestic), peaches (Prunus persica) and cherries (Prunus avium) remain labor-intensive Kaur et al. (2023); Rutledge and Mérel (2023). For example, Cherries require careful hand-picking due to their delicate nature, to prevent damage to both the fruit and the trees—a task that current machinery cannot perform with the necessary gentleness Karkee et al. (2018); Zhou et al. (2016). Apples and many other fruit crops are also harvested manually despite recent commercialization efforts Musacchi and Serra (2018). Plums (Prunus domestica) and apples need selective pruning to maintain desired canopy structure that is essential for improved fruit quality and manage tree health Seehuber et al. (2011). Furthermore, these crops depend heavily on manual thinning to optimally distribute fruit across canopy areas, which is essential to optimize fruit size and quality, and to minimize biennial bearing Iglesias et al. (2023); Costa and Botton (2022). These manual operations create a substantial reliance on decreasing human labor, highlighting the need for innovative labor-saving solutions.
@@
Among these crop-load management operations, immature green fruit thinning is one of the crucial operations used to optimize fruit size and quality by reducing competition for water, sunlight, and nutrients among fruits Dennis Jr (2000); Costa et al. (2018). This practice is particularly crucial for apple production, an important crop in the United States Statista (2022) and globally O’Rourke (2021). The labor-intensive nature of manual thinning (1), combined with persistent labor shortages and the increasing costs associated with securing skilled labor, highlight the urgent need for innovative solutions to automate this process Assirelli et al. (2018); Lordan et al. (2018). Each year, over a hundred thousand workers come to the US for farm labor Holmes and Ramirez-Lopez (2023); Weiler (2022). Figure 1 shows how these workers are engaged in physically demanding agricultural tasks such as green fruitlet thinning in commercial orchards.
@@
Figure 1: Manual thinning of apple fruitlet in a commercial ’Scilate’ apple orchard in Prosser, Washington State, USA, during 2024 growing season; (a) use of an aluminum ladder by farm workers to access fruit clusters in upper canopy region, demonstrating the physical demands and safety risks of traditional methods; b) manual thinning using a mechanized platform, which reduced the need for climbing ladders, but still required extensive manual efforts to reach fruitlet clusters; c) Showing the high-density examples of immature green apples in commercial orchards which needs thinning.
As shown in 1a, workers often move between trees with carrying and setting an aluminum ladder, and then perform these farming tasks while claiming up and down the ladder that are often 14 feet or higher. This method not only demands physical agility to climb and reposition the ladder continuously but also poses safety risks inherent in working at heights and with repeated hand movements.
Figure 1b depicts an alternative approach where multiple workers are using a mechanized platform that moves through orchard rows (either automated or driven by an operator). While this method reduces the need to climb and improves the working condition compared to ladder-based operation, it still requires significant manual effort. Workers must bend and stretch extensively to reach fruitlets in dense clusters, as shown in Figure 1c. This physically demanding work not only entails high labor costs but also poses significant health risks to workers, as mentioned before, particularly related to spinal injuries and other musculoskeletal disorders. Studies have highlighted that the repetitive and strenuous postures required in manual thinning, such as excessive back bending and upper arm elevation, can lead to severe physical strain Thamsuwan et al. (2020). Specifically, while platforms can reduce upper arm strain compared to ladders, they do not significantly alleviate the stress on workers’ backs, who must still engage in substantial torso flexion. Furthermore, the agricultural sector is notorious for the prevalence of upper limb injuries, with a systematic review revealing that a substantial number of these injuries occur during manual tasks like harvesting and thinning, often leading to long-term disabilities and economic consequences Mucci et al. (2020).
Despite some advancements in mechanization, manual labor remains prevalent in fruitlet thinning due to the precision required for tasks like immature apple detection, where occlusion by leaves, overlapping fruits, and variable lighting conditions present significant challenges Linker et al. (2012); Sun et al. (2019). Studies have explored deep learning approaches for detecting young apples with some success, yet issues with accuracy and processing speed persist Xia et al. (2018); Tian et al. (2019); Huang et al. (2021). Techniques like Lidar-based modeling for 3D pose estimation of green pepper where the orientation of a fruit in space is obtained by fitting a model to surface points of the fruit Eizentals and Oka (2016), low-cost RGB-D sensors with multiple three-dimensional (3D) line-segments detection Lin et al. (2019), and binocular imagery Yin et al. (2021) have shown high precision and recall in pose estimation for various fruits. However, these methods often struggle with real-time processing requirements and complex environments, making them less feasible for field applications. Therefore, there is a need for faster, more precise, and simpler fruit thinning systems capable of operating effectively in complex and dynamic orchard environments. While innovative approaches including the use of single RGB images for citrus pose estimation Sun et al. (2023) and detection of symmetry axes in 3D point clouds for in citrus pose estimation from an RGB image for automated harvestingin, and sweet peppers pose estimationLi et al. (2018) are promising, their dependence on specific conditions.
To address these challenges, this study focused on developing a machine vision system for 3D pose estimation of immature green fruits in commercial apple orchards, which will provide a foundation for the development of robotic fruit thinning technologies. The specific objectives are as follows
  • 1.
    Data Acquisition: Deploy a machine vision system on a robotic platform to collect high-resolution RGB images within a commercial orchard, creating a comprehensive dataset for model training and testing.
  • 2.
    Deep Learning-based Detection and Pose Estimation: Utilize the YOLO11 model to accurately detect and determine the position and orientation of immature green fruits, enhancing the precision with uniform hyperparameter settings across the models.
  • 3.
    RGB to RGB-D Mapping Using Vision Transformers: Convert RGB images into 3D point clouds using Vision Transformers, specifically the Dense Prediction Transformer (DPT) and Depth Anything V2, to improve pose estimations.
  • 4.
    Field Level Validation: Validate these estimations against field-level measurements of the major axis values of immature green apple fruits, ensuring the accuracy and reliability of the vision system for practical applications.
2 Related Work
@@ 2.1 Deep Learning-based Fruit Detection and Pose Estimation:
Traditional methods for apple and other fruit detection have mostly focused on mature fruit suitable for robotic harvesting and crop-load estimation, utilizing basic image processing as well as deep learning techniques. These methods, however, often face challenges in detecting immature apples (desired for robotic fruitlet thinning) in complex natural environments where young fruits exhibit colors and textures similar to foliage Kataoka et al. (1998); Xuan et al. (2020). To address this challenge, Xia et al. (2018) Xia et al. (2018) employed an innovative method combining adaptive green and blue chromatic aberration mapping with an iterative threshold segmentation algorithm, achieving a true positive rate of 88.51% and an F1-Measure of 90.29% in detecting young green fruit, which was a notable improvement compared to past studies in outdoor orchard environments.
Recent advancements in machine learning and deep learning have introduced more sophisticated models, such as YOLO series of models, which significantly enhance detection capabilities by learning from diverse datasets without explicit programming Liakos et al. (2018); Xuan et al. (2020). Utilizing these advancements, a few studies have focused on detecting immature green apples in complex orchard environments, which is crucial for automating the thinning process. For example, Tian et al. (2019) Tian et al. (2019) refined the YOLO-V3 model using DenseNet processing, enhancing apple detection capabilities across various growth stages under challenging conditions such as fluctuating illumination and fruit overlap, demonstrating the potential for real-time applications. Additionally, Huang et al. (2021) Huang et al. (2021) further advanced the YOLOv3 framework for identifying immature apples, achieving high-speed detection of 83 frames per second on a 1080ti GPU, effectively addressing occlusions and overlapping fruits. Despite these efforts, there is still a lack of studies targeting the development of comprehensive vision models capable of accurately estimating fruit location, orientation, and size, which is essential for developing robotic systems of immature green apple thinning. This task is particularly challenging because of the early growth stage of fruit, when issues such as occlusion and overlap of fruits by leaves, branches, and other fruit, as well as variable lighting conditions, are more prominent Linker et al. (2012); Sun et al. (2019).
It is also noted that there has been a wide range of studies on other fruit crops showing considerable advancements in fruit detection and pose estimation, albeit with shared limitations. Eizentals et al. (2016) Eizentals and Oka (2016) utilized Lidar-based model matching for 3D pose estimation of green peppers, achieving effective results under controlled conditions but struggling in natural, cluttered settings. Lin et al. (2019) Lin et al. (2019) applied a low-cost RGB-D sensor for guava detection, achieving high accuracy but facing challenges in real-time processing, highlighting the need for faster methods. Similarly, Yin et al. (2021) Yin et al. (2021) used binocular imagery and neural networks for detecting grape clusters, showing high accuracy but requiring complex setups that hinder practical field application. These studies underscore the potential of advanced imaging and machine learning in agricultural applications while also highlighting critical areas for improvement, particularly in processing speed and operational simplicity.
Studies in pose estimation have also extended to crops like tomatoes and peppers, with researchers addressing unique challenges in these domains. Kim et al. Kim et al. (2023) developed a 2D pose estimation method that effectively identified tomato-pedicel pairs, though its reliance on visible keypoints limited its effectiveness under occlusion. Li et al. (2018) Li et al. (2018) introduced a pose estimation strategy for sweet peppers using symmetry axes in point clouds, which varied in performance based on the physical state of the peduncle, introducing inconsistencies in uncontrolled environments. While the studies based on deep learning methods in fruit detection and pose estimation are expanding, the common challenges across the studies include the need for higher accuracy, faster processing speeds, and simpler, more robust systems capable of functioning effectively in dynamically changing and visually complex orchard environments.
@@ 2.2 Vision Transformers Application in Agriculture
Vision Transformers (ViTs), which utilize the self-attention mechanism to process images as sequences of patches, have emerged as powerful alternatives to convolutional neural networks for various image analysis tasks Thakur et al. (2021); Li et al. (2022) . ViTs have increasingly been recognized for their powerful capabilities in addressing complex challenges in agriculture, particularly in the realm of plant disease classification. Studies such as that by Salamai et al. (2023) Salamai et al. (2023) have showcased the effectiveness of ViTs in detecting paddy diseases, where the lesion-aware visual transformer model achieved an impressive accuracy of approximately 98.74% and an F1-score of around 98.18%. This performance highlights the potential of ViTs in precision agriculture for enhancing disease identification with high accuracy and efficiency. Similarly, Thai et al. (2023) Thai et al. (2023) introduced FormerLeaf, a transformer-based model that incorporates the Least Important Attention Pruning (LeIAP) technique and sparse matrix-matrix multiplication (SPMM) to optimize the detection of leaf diseases. This model not only reduced training times but also significantly improved accuracy, demonstrating the adaptability of ViTs in agricultural settings. Furthermore, Thakur et al. (2021) Thakur et al. (2021) combined conventional Convolutional Neural Networks (CNNs) with Vision Transformers to create PlantViT, a hybrid model for diagnosing plant diseases from leaf images, achieving notable accuracy across extensive datasets. This hybrid approach underscores the robustness and generalizability of ViTs in diverse agricultural environments.
@@
In addition to disease detection, ViTs have also been applied to other agricultural tasks such as crop yield prediction and multi-object tracking, illustrating their versatility. Lin et al. (2023) Lin et al. (2023) developed a Multi-Modal Spatial-Temporal Vision Transformer (MMST-ViT) model, which integrates visual remote sensing data with meteorological inputs to evaluate the influence of environmental factors on crop yields. This model utilizes a novel multi-modal contrastive learning technique for effective pre-training without extensive human intervention, significantly outperforming existing models in several key performance metrics. Specifically, MMST-ViT excelled in metrics such as Root Mean Square Error (RMSE), R-squared, and Pearson Correlation Coefficient, underscoring its superior predictive accuracy in agricultural settings. Additionally, Hernandez et al. (2024) Hernandez and Medeiros (2024) proposed a Multi-Object Tracking (MOT) framework using a Vision Transformer, tailored to agricultural environments. This framework employed the local feature matching transformer (LoFTR) to improve spatial associations among plants within cluttered fields, significantly enhancing the robustness and accuracy of robotic tracking applications. Tested on specialized datasets such as LettuceMOT and AppleMOT, this method achieved up to a 44% improvement in tracking accuracy over existing solutions Hernandez and Medeiros (2024). These advancements demonstrated the potential of ViTs to manage spatial dynamics effectively, facilitating the adoption and advancement of robotic technologies in agriculture.
While Vision Transformers have not yet been widely applied in automating orchard operations, their demonstrated success in various other agricultural applications shows the potential for exploring their applications to this area. Specifically, based on their robust performance in disease classification, crop yield prediction, and object tracking, there is a promising opportunity to utilize ViTs for enhancing object detection, segmentation, localization, and tracking in fruit tree canopies, thus helping improve automation and robotics technologies for orchard operations.